# Skewed score: a statistical framework to assess autograders

### v2

@@ Introduction
In the following sections, we begin by presenting the general framework. We then walk through a series of common questions to demonstrate how the framework can be used to answer them. We begin by explaining how autograder scores can be compared to human scores (Question 1.1), and how this comparison can be integrated into an LLM evaluation analysis (Question 1.2). We then gradually increase the complexity of the setup and show how different questions about autograder quality can be addressed within this integrated framework. Specifically, we show how to assess whether the graders are biased towards certain models being evaluated (Question 2) and how to quantify individual differences in the case of multiple graders (Question 3). We also show how to analyse item-level patterns (Question 4), which allows us to compute inter-rater agreement metrics with quantification of uncertainty and, crucially, to identify whether disagreement stems from noise or systematic bias (something that traditional metrics cannot disentangle). Finally, we show how this framework can be applied to pairwise judgments settings, in which LLM outputs are compared directly to one another (as opposed to being assigned absolute scores), how to quantify intransitive (e.g., cyclic) preferences and how to assess whether graders have biases towards longer formats (Question 5). In table 1, we provide a summary of the evaluation questions along with their corresponding formalisations as a practical guide for researchers using our framework.
@@
Is my grading scale well calibrated? Inspect cutpoints from ordered regression to analyse spacing and interpretability of score intervals. appendix C
@@
Imagine a typical scenario: a researcher, Florence [^1] In tribute to the pioneering work of two Florence Nightingales in statistics: the 19th-century nurse who applied statistical methods to public health and the 20th-century statistician who advanced combinatorics and statistical theory., is using LLMs to answer open-ended questions. To evaluate the quality of the model’s responses, she decides to manually grade them. She develops a detailed rubric (table 2) with scores ranging from 1 (completely off-topic) to 10 (demonstrating deep understanding and insightful connections).
@@
To illustrate how the framework can be applied in practice, we will follow Florence as she navigates common challenges in evaluating LLM outputs with autograders. All examples are based on simulated data, which can be reproduced from a publicly available repository. To facilitate wider adoption, all GLMs presented in this paper are implemented in the open-source HiBayes package. table 1 summarises the evaluation questions addressed, their corresponding GLM formulation, and the sections where each is presented and explained in detail.
@@ 1.1 Quantify the mean difference
As Florence wants to validate the autograder before using it, she decides to run it on the same N=100 examples that she had graded, and compare the results. In the left panel of fig. 1, we see a violin plot of the simulated scores given by Florence (human expert) in green, and by the autograder in yellow. Using our GLM framework we can fit a regression model with parameter estimates that quantifies the effect of the grader on the assigned score (eq. 2). Because the score can take values [1-10], we use an ordered logistic likelihood function, which accounts for the finite range and the categorical nature of the outcomes. This ordered logistic model has two main components: the linear predictor (cf. eq. 1) and the cutpoints. In our case, the linear predictor produces a continuous value on a latent (unobserved) scale. The ordered logistic model assumes that our discrete 1-10 scores are actually observations of this underlying continuous scale, divided by cutpoints. When the linear predictor falls between particular cutpoints $c_{j}$, we observe the corresponding discrete score. For example, if the linear predictor value falls between cutpoints $c_{3}$ and $c_{4}$, we would observe a score of 4. These cutpoints are estimated from the data along with the $\beta$ coefficients (see appendix C for more technical details).
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into discrete score categories, and the ordered logistic model uses a cumulative logit link function to map from the linear predictor to score probabilities through the relation: $\text{logit}(P(\text{score}_{i}\leq j))=c_{j}-\phi_{i}.$ The variable $X_{i}^{\text{grader}}$ encodes the identity of the grader who assigned score $i$ using effect coding. [^2] Effect coding is a contrast coding scheme used in to represent categorical variables. Each level of the variable is assigned a numerical value such that the resulting coded variables are centered and sum to zero across all groups. This centering allows the intercept to reflect the grand mean (the average of the group means), and the coefficients represent deviations from that mean. The linear model has two parameters: an intercept $\beta_{0}$, and a grader effect $\beta_{1}$. The intercept $\beta_{0}$ represents the average latent score across both graders, due to the centering imposed by effect coding. The grader coefficient $\beta_{1}$ captures the effect of the grader on the latent score and quantifies half the difference between the autograder and the human grader. Since the model encodes the autograder as $+1$ and the human as $-1$, the linear predictor shifts depending on the grader identity. These parameters define the location of the latent score, which is then mapped to the observed 1-10 categories via the threshold cutpoints. This shift in the latent score translates to a systematic difference in the probabilities of assigning higher or lower scores.
@@
To evaluate hypotheses, Florence fits different models to the data and compares their relative fit using statistical criteria such as Leave-One-Out cross-validation (LOO-CV) or the Widely Applicable Information Criterion (WAIC). For instance, here, Florence should compare a model that includes a grader effect (eq. 2) against a simpler model without this effect (i.e., $\phi_{i}=\beta_{0}$). The former statistical model tests the hypothesis that grader identity affects the assigned score, while the latter one serves as a null model with no grader effect. Model comparison indicates whether adding the grader effect improves model fit, but it does not guarantee that the grader parameter itself is significant. Therefore, once the best-fitting model (often called “winning model”) is identified, Florence should inspect the parameter estimates and their credible intervals to confirm if there is clear statistical evidence for the grader effect.
@@
After fitting the statistical models and running a model comparison (fig. 8 in appendix B), we identify the model with the grader effect (eq. 2) as the best-fitting model. We can then examine the estimated effect of interest of this model (right panel of fig. 1). Since the grader variable is effect coded as $+1$ for the autograder and $-1$ for the human, the plotted difference in latent score between the autograder and the human grader is $2\beta_{1}$. This difference is defined as autograder minus human, so a negative value indicates that the human grader assigns higher scores on the latent scale. We observe a negative difference, with credible intervals that do not include zero, indicating a meaningful difference between the two graders. This confirms that the autograder systematically assigns lower scores than the human grader.
@@
If the observed difference in scores seems small, Florence might also want to assess whether this difference is negligible for practical purposes. One way to do this is by defining a Region of Practical Equivalence (ROPE; [Kruschke, 2018]), an interval around zero within which effect sizes are considered too small to matter. A commonly used ROPE for standardized effect sizes is from $-0.1$ to $0.1$, which corresponds to approximately $-0.18$ to $0.18$ on the log-odds scale used by logistic models. If the entire credible interval of the grader coefficient falls within this ROPE, Florence can conclude that the autograder and the human grader are practically equivalent in terms of their scoring.
@@ 1.2 Integrating the autograder assessment with the research question
Now suppose that the $N=100$ graded responses (each graded by both Florence and her autograder) are in fact 50 unique items, with each item answered once by LLM A and once by LLM B. The data might look like the left panel of fig. 2. To answer the two questions above, we can fit a regression model with an intercept $\beta_{0}$, which represents the overall average latent score, and a coefficient $\beta_{1}$ that quantifies the effect of the grader, as before (eq. 2). In addition, we include a coefficient $\beta_{2}$ to quantify the effect of LLM A versus LLM B on the score.
@@
The variables $X_{i}^{\text{grader}}$ and $X_{i}^{\text{LLM}}$ encode, respectively, the identity of the grader and the LLM associated with score $i$, using effect coding. Each variable takes a value of $+1$ or $-1$ to distinguish between the two levels (e.g., autograder vs. human, LLM A vs. LLM B). This coding imposes a sum-to-zero constraint on each effect, so that $\beta_{1}$ and $\beta_{2}$ represent half the difference in latent score between levels of the corresponding variable.
The coefficients $\beta_{1}$ and $\beta_{2}$ allow us to answer the two questions above in a single statistical model. After fitting the model, we can make inferences based on the effect sizes represented by the coefficients (right panel of fig. 2). We observe two things:
@@
Going back to Florence, she recently learned about potential self-bias of autograders and is concerned that her autograder (from model family A) might unfairly favour outputs from LLM A (also from model family A). To assess whether such self-bias exists, she uses a second autograder (from model family B). She wants to investigate whether responses from LLM A receive higher scores when graded by the autograder A compared to when graded by autograder B (and vice versa). The resulting data are shown in the left panel of fig. 3.
To answer the question about self-bias, we extend the previous model (eq. 3) by adding a term that captures whether specific graders systematically favour outputs from specific LLMs. This is implemented as a set of grader–LLM interaction effects [^3] Strictly speaking, this is not a single interaction effect (e.g., $\beta_{3}\cdot X^{\text{grader}}\cdot X^{\text{LLM}}$), but a set of parameters estimated independently using index-based coding (where each grader–LLM combination has a unique integer index). This allows direct comparisons across specific combinations rather than relying on a single coefficient., denoted by $\beta_{3}$.
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is a vector of cutpoints, and $X_{i}^{\text{grader}}$ and $X_{i}^{\text{LLM}}$ identify the grader and LLM assigned to observation $i$. As we now have more than two graders, the main effect $\beta_{1}$ is a vector of coefficients representing each grader’s deviation from the grand mean score, estimated using effect coding. The LLM variable still has only two levels and is binary-coded as before (see eq. 3), so $\beta_{2}$ is a scalar coefficient applied to $X_{i}^{\text{LLM}}\in\{-1,+1\}$. The interaction term $\beta_{3}$ is a matrix of parameters estimated using index-based coding, with one distinct coefficient independently estimated for each grader–LLM combination. Unlike effect coding, which estimates deviations from a grand mean (i.e., the average across all grader–LLM combinations), and dummy coding, which estimates differences from a reference category, index-based coding directly estimates each combination independently relative to the intercept. This allows for direct interpretation of each pair (e.g., how Autograder scores LLM A compared to the intercept) and simplifies comparisons across combinations.
After selecting the best-fitting model using model comparison techniques (fig. 9 in appendix B), we examine the estimated effects (right panel of fig. 3). The interaction parameters $\beta_{3}$ include a positive effect for the combination of Autograder A and LLM A, indicating that Autograder A tends to assign higher scores to LLM A relative to the overall baseline (i.e., the intercept). Conversely, the interaction term for Autograder B and LLM A is negative, suggesting that Autograder B gives lower scores to LLM A compared to the same baseline. While these individual patterns show grader-specific scoring tendencies, the key evidence for self-bias emerges from directly contrasting how Autograder A scores LLM A against how Autograder A scores LLM B. Looking at these specific effects, we see a difference in their posterior means suggesting a tendency toward self-bias (although, because their credible intervals partially overlap, additional data would be necessary for a more definitive conclusion).
@@
In practice, Florence will probably not have the time (nor the motivation) to grade even N=100 responses herself. Instead, she will ask a few of her colleagues to help grade some of the responses. Instead of only having grades from one human grader, she will now have grades from multiple humans, and can and can assess their individual scoring patterns (Human X, Y and Z in the left panel of fig. 4.) Additionally, because she is aware of potential autograder biases, she might want to try other autograders, for instance LLMs from different model families, or LLMs from the same family but using different prompts. Just like with the humans, she will now have grades from different autograders that she can examine individually (Autograder A, B and C in the left panel of fig. 4.).
In terms of analysis, she could account for each grader individually, similarly to the $\beta_{1}$ effect in eq. 3 (visualised in the right panel of fig. 3). The problem with that method is that the number of parameters increases with the number of graders, potentially complicating interpretation.
@@
To capture such a group-level difference, we can define what is known as a hierarchical GLM (cf. eq. 5). In this model, each individual grader has their own average scoring tendency ($\beta_{\text{grader}_{i}}$), but these grader-specific effects are drawn from a group-level distribution depending on their type (human or autograder). This approach allows us to estimate separate group-level means for human graders and autograders, which can then be directly compared to assess systematic differences between the two groups. It also captures individual-level deviations, which reflect how each grader differs from the average behaviour of their group. This makes it possible to detect outliers or particularly lenient or strict graders. By sharing information across graders of the same type, a property known as partial pooling, the statistical model makes more efficient use of limited data, which is particularly helpful when some graders have relatively few observations.
@@
where $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is a vector of cutpoints that divide the latent scale into discrete score categories. The intercept $\beta_{0}$ represents the grand mean, and $\beta_{2}$ is a scalar coefficient applied to $X_{i}^{\text{LLM}}\in\{-1,+1\}$, which indicates whether the response was generated by LLM A or LLM B. Here, $\beta_{1}(X_{i}^{\text{grader}})$, unlike before, represents the effect of the individual grader who assigned score $i$, and is drawn from a group-level distribution based on grader type (human or autograder). Specifically, $\beta_{1}(X_{i}^{\text{grader}})\sim\mathcal{N}(\mu_{\text{graderType}_{i}},\sigma^{2}_{\text{graderType}_{i}})$, where $\mu_{\text{graderType}_{i}}$ represents the average score tendency for each grader type, and $\sigma^{2}_{\text{graderType}_{i}}$ captures variability within each type. The prior distributions for the group-level means and variances ($\mu_{\text{graderType}}$ and $\sigma^{2}_{\text{graderType}}$) are specified in appendix A. This hierarchical structure enables the model to estimate both the average difference between human and autograder scores and the variation among individual graders.
To formally assess whether this group-level difference exists, Florence should of course compare this hierarchical approach against the simpler flat model (cf. fig. 10 in appendix B for a model comparison). Here we select the hierarchical model to demonstrate how to examine both group-level differences between grader types and individual grader characteristics.
In the right panel of fig. 4 we see the individual grader effects ($\beta_{1}$; Autograder A–C and Human X–Z in the plot) that provide information about how each grader scores on average. The group-level means for human and autograders ($\mu_{\text{graderType}}$; human mean and Autograder mean in the plot) allow us to estimate the average difference in scoring tendencies between the two groups. The results show that human graders systematically assign higher scores than autograders, regardless of individual differences.
@@
Figure 5: Illustration of how to use Bayesian GLMs to address Question 4 (How do scores differ at an item level?) using simulated data. Left panel: Violin plot of simulated scores for each item (1–4), grouped by LLM and grader identity. Each cell shows the distribution of scores assigned by a given grader to responses from a particular model on a given item. Right panel: Posterior distributions of estimated effects from the item-level GLM (eq. 6). The plot shows main effects for grader and LLM identity (top), item main effects (bottom), and grader–item interactions (middle). Horizontal blue lines represent 95% credible intervals, and the dashed red vertical line indicates a null effect ($\beta=0$). Item 1 has a strong positive effect, suggesting it consistently receives higher scores. In contrast, Item 4 has a negative effect, indicating that it receives lower scores. Grader–item interaction terms are small and uncertain, indicating no evidence of systematic grader disagreement on specific items.
@@
To answer these questions, she needs repeated responses for the same items. Until now, we have assumed that each data point corresponds to a different item, so lets instead imagine that Florence’s dataset consists of four items (for simplicity), with each model answering each item 25 times. This results in 50 responses per item (25 per model); each grader scores all of the attempts. The data split by items can be seen in the left panel of fig. 5 (different items are represented by violin plots of the same colour).
To answer Question 4, we extend eq. 3 by including two additional terms. The first term, $\beta_{3}(X_{i}^{\text{item}})$, accounts for a main effect of items, capturing whether some items receive systematically higher or lower scores. The second term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, represents a grader–item interaction, allowing us to test whether particular graders behave differently on specific items.
@@
Here, $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into ordered categories. The term $\beta_{1}(X_{i}^{\text{grader}})$ captures the main effect of each grader, and $\beta_{2}$ models the effect of LLM identity (e.g., whether the response was produced by LLM A or B). As mentioned above, the new term $\beta_{3}(X_{i}^{\text{item}})$ represents the main effect of each item, which captures whether some questions tend to receive higher or lower scores overall. The final term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, captures grader–item interactions and is implemented in the same way as the interaction term in eq. 4: rather than a single multiplicative term, this is a set of coefficients (one for each grader–item combination) estimated using dummy coding without a sum-to-zero constraint. This allows us to directly compare individual combinations and detect whether certain graders are more lenient or harsh on specific items. As before, all main categorical effects (grader, item) are encoded using effect coding, so that the resulting coefficients reflect deviations from the overall mean.
After fitting the model, Florence inspects the estimated effects (right panel of fig. 5). She focuses on two aspects:
@@
    The main effect of item ($\beta_{3}$ in eq. 6), observing that some items consistently receive higher or lower scores. For example, Item 1 shows a notably positive main effect, indicating that graders tend to assign higher scores to responses for this item. Conversely, Item 4 exhibits a negative effect, implying that graders give lower scores on this item. Overall, this suggests that Item 1 is easier to answer, while Item 4 is more challenging.
@@
    The grader–item interaction term ($\beta_{4}$ in eq. 6), which helps identify whether graders differ more on specific questions. Most interaction estimates are small and their credible intervals overlap with zero, indicating weak evidence that specific graders behave differently on individual items.

### v3

@@ Introduction
In the following sections, we begin by presenting the general framework. We then walk through a series of common questions to demonstrate how the framework can be used to answer them. We begin by explaining how autograder scores can be compared to human scores (Question 1.1), and how this comparison can be integrated into an LLM evaluation analysis (Question 1.2). We then gradually increase the complexity of the setup and show how different questions about autograder quality can be addressed within this integrated framework. Specifically, we show how to assess whether the graders are biased towards certain models being evaluated (Question 2) and how to quantify individual differences in the case of multiple graders (Question 3). We also show how to analyse item-level patterns (Question 4), which allows us to compute inter-rater agreement metrics with quantification of uncertainty and, crucially, to identify whether disagreement stems from noise or systematic bias (something that traditional metrics cannot disentangle). Finally, we show how this framework can be applied to pairwise judgments settings, in which LLM outputs are compared directly to one another (as opposed to being assigned absolute scores), how to quantify intransitive (e.g., cyclic) preferences and how to assess whether graders have biases towards longer formats (Question 5). In table˜1, we provide a summary of the evaluation questions along with their corresponding formalisations as a practical guide for researchers using our framework.
@@
Is my grading scale well calibrated? Inspect cutpoints from ordered regression to analyse spacing and interpretability of score intervals. appendix˜C
@@
Imagine a typical scenario: a researcher, Florence [^1] In tribute to the pioneering work of two Florence Nightingales in statistics: the 19th-century nurse who applied statistical methods to public health and the 20th-century statistician who advanced combinatorics and statistical theory., is using LLMs to answer open-ended questions. To evaluate the quality of the model’s responses, she decides to manually grade them. She develops a detailed rubric (table˜2) with scores ranging from 1 (completely off-topic) to 10 (demonstrating deep understanding and insightful connections).
@@
To illustrate how the framework can be applied in practice, we will follow Florence as she navigates common challenges in evaluating LLM outputs with autograders. All examples are based on simulated data, which can be reproduced from a publicly available repository. To facilitate wider adoption, all GLMs presented in this paper are implemented in the open-source HiBayes package. table˜1 summarises the evaluation questions addressed, their corresponding GLM formulation, and the sections where each is presented and explained in detail.
@@ 1.1 Quantify the mean difference
As Florence wants to validate the autograder before using it, she decides to run it on the same N=100 examples that she had graded, and compare the results. In the left panel of fig.˜1, we see a violin plot of the simulated scores given by Florence (human expert) in green, and by the autograder in yellow. Using our GLM framework we can fit a regression model with parameter estimates that quantifies the effect of the grader on the assigned score (eq.˜2). Because the score can take values [1-10], we use an ordered logistic likelihood function, which accounts for the finite range and the categorical nature of the outcomes. This ordered logistic model has two main components: the linear predictor (cf. eq.˜1) and the cutpoints. In our case, the linear predictor produces a continuous value on a latent (unobserved) scale. The ordered logistic model assumes that our discrete 1-10 scores are actually observations of this underlying continuous scale, divided by cutpoints. When the linear predictor falls between particular cutpoints $c_{j}$, we observe the corresponding discrete score. For example, if the linear predictor value falls between cutpoints $c_{3}$ and $c_{4}$, we would observe a score of 4. These cutpoints are estimated from the data along with the $\beta$ coefficients (see appendix˜C for more technical details).
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into discrete score categories, and the ordered logistic model uses a cumulative logit link function to map from the linear predictor to score probabilities through the relation: $\text{logit}(P(\text{score}_{i}\leq j))=c_{j}-\phi_{i}.$ The variable $X_{i}^{\text{grader}}$ encodes the identity of the grader who assigned score $i$ using effect coding. [^2] Effect coding is a contrast coding scheme used in to represent categorical variables. Each level of the variable is assigned a numerical value such that the resulting coded variables are centered and sum to zero across all groups. This centering allows the intercept to reflect the grand mean (the average of the group means), and the coefficients represent deviations from that mean. The linear model has two parameters: an intercept $\beta_{0}$, and a grader effect $\beta_{1}$. The intercept $\beta_{0}$ represents the average latent score across both graders, due to the centering imposed by effect coding. The grader coefficient $\beta_{1}$ captures the effect of the grader on the latent score and quantifies half the difference between the autograder and the human grader. Since the model encodes the autograder as +1+1 and the human as 1-1, the linear predictor shifts depending on the grader identity. These parameters define the location of the latent score, which is then mapped to the observed 1-10 categories via the threshold cutpoints. This shift in the latent score translates to a systematic difference in the probabilities of assigning higher or lower scores.
@@
To evaluate hypotheses, Florence fits different models to the data and compares their relative fit using statistical criteria such as Leave-One-Out cross-validation (LOO-CV) or the Widely Applicable Information Criterion (WAIC). For instance, here, Florence should compare a model that includes a grader effect (eq.˜2) against a simpler model without this effect (i.e., $\phi_{i}=\beta_{0}$). The former statistical model tests the hypothesis that grader identity affects the assigned score, while the latter one serves as a null model with no grader effect. Model comparison indicates whether adding the grader effect improves model fit, but it does not guarantee that the grader parameter itself is significant. Therefore, once the best-fitting model (often called “winning model”) is identified, Florence should inspect the parameter estimates and their credible intervals to confirm if there is clear statistical evidence for the grader effect.
@@
After fitting the statistical models and running a model comparison (fig.˜8 in appendix˜B), we identify the model with the grader effect (eq.˜2) as the best-fitting model. We can then examine the estimated effect of interest of this model (right panel of fig.˜1). Since the grader variable is effect coded as +1+1 for the autograder and 1-1 for the human, the plotted difference in latent score between the autograder and the human grader is $2\beta_{1}$. This difference is defined as autograder minus human, so a negative value indicates that the human grader assigns higher scores on the latent scale. We observe a negative difference, with credible intervals that do not include zero, indicating a meaningful difference between the two graders. This confirms that the autograder systematically assigns lower scores than the human grader.
@@
If the observed difference in scores seems small, Florence might also want to assess whether this difference is negligible for practical purposes. One way to do this is by defining a Region of Practical Equivalence (ROPE; [Kruschke, 2018]), an interval around zero within which effect sizes are considered too small to matter. A commonly used ROPE for standardized effect sizes is from 0.1-0.1 to $0.1$, which corresponds to approximately 0.18-0.18 to $0.18$ on the log-odds scale used by logistic models. If the entire credible interval of the grader coefficient falls within this ROPE, Florence can conclude that the autograder and the human grader are practically equivalent in terms of their scoring.
@@ 1.2 Integrating the autograder assessment with the research question
Now suppose that the $N=100$ graded responses (each graded by both Florence and her autograder) are in fact 50 unique items, with each item answered once by LLM A and once by LLM B. The data might look like the left panel of fig.˜2. To answer the two questions above, we can fit a regression model with an intercept $\beta_{0}$, which represents the overall average latent score, and a coefficient $\beta_{1}$ that quantifies the effect of the grader, as before (eq.˜2). In addition, we include a coefficient $\beta_{2}$ to quantify the effect of LLM A versus LLM B on the score.
@@
The variables $X_{i}^{\text{grader}}$ and $X_{i}^{\text{LLM}}$ encode, respectively, the identity of the grader and the LLM associated with score $i$, using effect coding. Each variable takes a value of +1+1 or 1-1 to distinguish between the two levels (e.g., autograder vs. human, LLM A vs. LLM B). This coding imposes a sum-to-zero constraint on each effect, so that $\beta_{1}$ and $\beta_{2}$ represent half the difference in latent score between levels of the corresponding variable.
The coefficients $\beta_{1}$ and $\beta_{2}$ allow us to answer the two questions above in a single statistical model. After fitting the model, we can make inferences based on the effect sizes represented by the coefficients (right panel of fig.˜2). We observe two things:
@@
Going back to Florence, she recently learned about potential self-bias of autograders and is concerned that her autograder (from model family A) might unfairly favour outputs from LLM A (also from model family A). To assess whether such self-bias exists, she uses a second autograder (from model family B). She wants to investigate whether responses from LLM A receive higher scores when graded by the autograder A compared to when graded by autograder B (and vice versa). The resulting data are shown in the left panel of fig.˜3.
To answer the question about self-bias, we extend the previous model (eq.˜3) by adding a term that captures whether specific graders systematically favour outputs from specific LLMs. This is implemented as a set of grader–LLM interaction effects [^3] Strictly speaking, this is not a single interaction effect (e.g., $\beta_{3}\cdot X^{\text{grader}}\cdot X^{\text{LLM}}$), but a set of parameters estimated independently using index-based coding (where each grader–LLM combination has a unique integer index). This allows direct comparisons across specific combinations rather than relying on a single coefficient., denoted by $\beta_{3}$.
@@
where $\phi_{i}$ is the linear predictor on the latent scale, $\boldsymbol{c}$ is a vector of cutpoints, and $X_{i}^{\text{grader}}$ and $X_{i}^{\text{LLM}}$ identify the grader and LLM assigned to observation $i$. As we now have more than two graders, the main effect $\beta_{1}$ is a vector of coefficients representing each grader’s deviation from the grand mean score, estimated using effect coding. The LLM variable still has only two levels and is binary-coded as before (see eq.˜3), so $\beta_{2}$ is a scalar coefficient applied to $X_{i}^{\text{LLM}}\in\{-1,+1\}$. The interaction term $\beta_{3}$ is a matrix of parameters estimated using index-based coding, with one distinct coefficient independently estimated for each grader–LLM combination. Unlike effect coding, which estimates deviations from a grand mean (i.e., the average across all grader–LLM combinations), and dummy coding, which estimates differences from a reference category, index-based coding directly estimates each combination independently relative to the intercept. This allows for direct interpretation of each pair (e.g., how Autograder scores LLM A compared to the intercept) and simplifies comparisons across combinations.
After selecting the best-fitting model using model comparison techniques (fig.˜9 in appendix˜B), we examine the estimated effects (right panel of fig.˜3). The interaction parameters $\beta_{3}$ include a positive effect for the combination of Autograder A and LLM A, indicating that Autograder A tends to assign higher scores to LLM A relative to the overall baseline (i.e., the intercept). Conversely, the interaction term for Autograder B and LLM A is negative, suggesting that Autograder B gives lower scores to LLM A compared to the same baseline. While these individual patterns show grader-specific scoring tendencies, the key evidence for self-bias emerges from directly contrasting how Autograder A scores LLM A against how Autograder A scores LLM B. Looking at these specific effects, we see a difference in their posterior means suggesting a tendency toward self-bias (although, because their credible intervals partially overlap, additional data would be necessary for a more definitive conclusion).
@@
In practice, Florence will probably not have the time (nor the motivation) to grade even N=100 responses herself. Instead, she will ask a few of her colleagues to help grade some of the responses. Instead of only having grades from one human grader, she will now have grades from multiple humans, and can and can assess their individual scoring patterns (Human X, Y and Z in the left panel of fig.˜4.) Additionally, because she is aware of potential autograder biases, she might want to try other autograders, for instance LLMs from different model families, or LLMs from the same family but using different prompts. Just like with the humans, she will now have grades from different autograders that she can examine individually (Autograder A, B and C in the left panel of fig.˜4.).
In terms of analysis, she could account for each grader individually, similarly to the $\beta_{1}$ effect in eq.˜3 (visualised in the right panel of fig.˜3). The problem with that method is that the number of parameters increases with the number of graders, potentially complicating interpretation.
@@
To capture such a group-level difference, we can define what is known as a hierarchical GLM (cf. eq.˜5). In this model, each individual grader has their own average scoring tendency ($\beta_{\text{grader}_{i}}$), but these grader-specific effects are drawn from a group-level distribution depending on their type (human or autograder). This approach allows us to estimate separate group-level means for human graders and autograders, which can then be directly compared to assess systematic differences between the two groups. It also captures individual-level deviations, which reflect how each grader differs from the average behaviour of their group. This makes it possible to detect outliers or particularly lenient or strict graders. By sharing information across graders of the same type, a property known as partial pooling, the statistical model makes more efficient use of limited data, which is particularly helpful when some graders have relatively few observations.
@@
where $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is a vector of cutpoints that divide the latent scale into discrete score categories. The intercept $\beta_{0}$ represents the grand mean, and $\beta_{2}$ is a scalar coefficient applied to $X_{i}^{\text{LLM}}\in\{-1,+1\}$, which indicates whether the response was generated by LLM A or LLM B. Here, $\beta_{1}(X_{i}^{\text{grader}})$, unlike before, represents the effect of the individual grader who assigned score $i$, and is drawn from a group-level distribution based on grader type (human or autograder). Specifically, $\beta_{1}(X_{i}^{\text{grader}})\sim\mathcal{N}(\mu_{\text{graderType}_{i}},\sigma^{2}_{\text{graderType}_{i}})$, where $\mu_{\text{graderType}_{i}}$ represents the average score tendency for each grader type, and $\sigma^{2}_{\text{graderType}_{i}}$ captures variability within each type. The prior distributions for the group-level means and variances ($\mu_{\text{graderType}}$ and $\sigma^{2}_{\text{graderType}}$) are specified in appendix˜A. This hierarchical structure enables the model to estimate both the average difference between human and autograder scores and the variation among individual graders.
To formally assess whether this group-level difference exists, Florence should of course compare this hierarchical approach against the simpler flat model (cf. fig.˜10 in appendix˜B for a model comparison). Here we select the hierarchical model to demonstrate how to examine both group-level differences between grader types and individual grader characteristics.
In the right panel of fig.˜4 we see the individual grader effects ($\beta_{1}$; Autograder A–C and Human X–Z in the plot) that provide information about how each grader scores on average. The group-level means for human and autograders ($\mu_{\text{graderType}}$; human mean and Autograder mean in the plot) allow us to estimate the average difference in scoring tendencies between the two groups. The results show that human graders systematically assign higher scores than autograders, regardless of individual differences.
@@
Figure 5: Illustration of how to use Bayesian GLMs to address Question 4 (How do scores differ at an item level?) using simulated data. Left panel: Violin plot of simulated scores for each item (1–4), grouped by LLM and grader identity. Each cell shows the distribution of scores assigned by a given grader to responses from a particular model on a given item. Right panel: Posterior distributions of estimated effects from the item-level GLM (eq.˜6). The plot shows main effects for grader and LLM identity (top), item main effects (bottom), and grader–item interactions (middle). Horizontal blue lines represent 95% credible intervals, and the dashed red vertical line indicates a null effect ($\beta=0$). Item 1 has a strong positive effect, suggesting it consistently receives higher scores. In contrast, Item 4 has a negative effect, indicating that it receives lower scores. Grader–item interaction terms are small and uncertain, indicating no evidence of systematic grader disagreement on specific items.
@@
To answer these questions, she needs repeated responses for the same items. Until now, we have assumed that each data point corresponds to a different item, so lets instead imagine that Florence’s dataset consists of four items (for simplicity), with each model answering each item 25 times. This results in 50 responses per item (25 per model); each grader scores all of the attempts. The data split by items can be seen in the left panel of fig.˜5 (different items are represented by violin plots of the same colour).
To answer Question 4, we extend eq.˜3 by including two additional terms. The first term, $\beta_{3}(X_{i}^{\text{item}})$, accounts for a main effect of items, capturing whether some items receive systematically higher or lower scores. The second term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, represents a grader–item interaction, allowing us to test whether particular graders behave differently on specific items.
@@
Here, $\phi_{i}$ is the linear predictor on the latent scale, and $\boldsymbol{c}$ is the vector of cutpoints that divide the latent scale into ordered categories. The term $\beta_{1}(X_{i}^{\text{grader}})$ captures the main effect of each grader, and $\beta_{2}$ models the effect of LLM identity (e.g., whether the response was produced by LLM A or B). As mentioned above, the new term $\beta_{3}(X_{i}^{\text{item}})$ represents the main effect of each item, which captures whether some questions tend to receive higher or lower scores overall. The final term, $\beta_{4}(X_{i}^{\text{grader}},X_{i}^{\text{item}})$, captures grader–item interactions and is implemented in the same way as the interaction term in eq.˜4: rather than a single multiplicative term, this is a set of coefficients (one for each grader–item combination) estimated using dummy coding without a sum-to-zero constraint. This allows us to directly compare individual combinations and detect whether certain graders are more lenient or harsh on specific items. As before, all main categorical effects (grader, item) are encoded using effect coding, so that the resulting coefficients reflect deviations from the overall mean.
After fitting the model, Florence inspects the estimated effects (right panel of fig.˜5). She focuses on two aspects:
@@
    The main effect of item ($\beta_{3}$ in eq.˜6), observing that some items consistently receive higher or lower scores. For example, Item 1 shows a notably positive main effect, indicating that graders tend to assign higher scores to responses for this item. Conversely, Item 4 exhibits a negative effect, implying that graders give lower scores on this item. Overall, this suggests that Item 1 is easier to answer, while Item 4 is more challenging.
@@
    The grader–item interaction term ($\beta_{4}$ in eq.˜6), which helps identify whether graders differ more on specific questions. Most interaction estimates are small and their credible intervals overlap with zero, indicating weak evidence that specific graders behave differently on individual items.